# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ Appendix C Dispersion of magnons
Appendix D Derivation of the one-magnon scattering amplitudes in the spin-orbit-free limit
In this Appendix we derive the result in Eq. (41) for the two-magnon-mediated scattering amplitudes in the $\theta_{0}=\pi/2$ limit, appropriate for vanishing spin-orbit interaction. Let us revisit Eq. (24) in the $\theta_{0}=\pi/2$ limit, where $Q^{\phi}_{s}=Q^{\theta}_{s}$. Using the eigenvector basis specified in Eq. (63) one finds magnon Green’s functions
$\mathcal{G}_{s\pm}(q)\sim(\mp i\gamma|\nu|+D{\boldsymbol{q}}^{2})^{-1}$ (70)
for the soft magnon branches. Projection onto those modes follows from
$\delta m^{\phi}_{\xi}(q)\rightarrow\frac{1}{\sqrt{2}}[\delta m_{s-}(q)+\delta m_{s+}(q)]~{},~{}\delta m^{\theta}_{\xi}(q)\rightarrow-\frac{i{\rm sgn}(\nu)}{\sqrt{2}}[\delta m_{s-}(q)-\delta m_{s+}(q)].$ (71)
The electron-magnon coupling accordingly evolves to
$\delta{\boldsymbol{m}}_{\xi}(q)\cdot{\boldsymbol{s}}_{\xi}(-q)\rightarrow\frac{1}{\sqrt{2}}\left\{\delta m_{s-}(q)[S^{y}_{\xi}(-q)-i{\rm sgn}(\nu)S^{x}_{\xi}(-q)]+\delta m_{s+}(q)[S^{y}_{\xi}(-q)+i{\rm sgn}(\nu)S^{x}_{\xi}(-q)]\right\},$ (72)
and after integrating out the magnons we obtain
$\mathcal{S}_{\rm I}=$ $-\frac{1}{4}\sum_{\xi_{1},\xi_{2}}\int_{q}\{\mathcal{G}_{s-}(q)[S^{y}_{\xi_{1}}(q)+i{\rm sgn}(\nu)S^{x}_{\xi_{1}}(q)][S^{y}_{\xi_{2}}(-q)-i{\rm sgn}(\nu)S^{x}_{\xi_{2}}(-q)]$
$+\mathcal{G}_{s+}(q)[S^{y}_{\xi_{1}}(q)-i{\rm sgn}(\nu)S^{x}_{\xi_{1}}(q)][S^{y}_{\xi_{2}}(-q)+i{\rm sgn}(\nu)S^{x}_{\xi_{2}}(-q)]\}.$ (73)
Comparing to Eq. (33) and expanding the Green’s functions for small $q$ yields the associated scattering amplitudes
$g^{-+,-+}_{\xi_{1},\xi_{2}}(q)=0,~{}~{}~{}~{}g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim-\frac{1}{-i\gamma\nu+D{\boldsymbol{q}}^{2}},$ (74)
as quoted in Eq. (41) in the main text.
Appendix E Derivation of two-magnon scattering amplitudes
In this Appendix we derive the two-magnon scattering amplitudes used in the main text. We first consider the generic case, and then separately treat the spin-orbit-free limit. We finally discuss how gapping out the Goldstone magnons (using a Zeeman field) regularizes the low-energy divergence of two-magnon scattering amplitudes.
E.1 Generic case
We first analyze the scattering amplitudes in the generic case where $\theta_{0}\neq 0,\pi/2$. As shown in Eqs. (44) and (45) of the main text, the pairing interaction arising from the ladder diagram $g_{\rm l}^{(2)}$ and from the crossed diagram $g_{\rm x}^{(2)}$ are asymptotically the same, since the asymptotic frequency-momentum-dependence of $g^{+-,-+}_{K,K^{\prime}}$ and $g^{+-,-+}_{K,K^{\prime}}$ at small $q$ are identical. The amplitudes of these two processes are given by
$g^{(2)}_{\rm l}(q)=g^{(2)}_{\rm x}(q)=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int_{\Omega,{\boldsymbol{Q}}}\frac{1}{(\Omega+\nu)^{2}+c_{s}^{2}({\boldsymbol{Q}}+{\boldsymbol{q}})^{2}}\frac{1}{\Omega^{2}+c_{s}^{2}{\boldsymbol{Q}}^{2}}.$ (75)
Here, we have used $\Omega\ll\bar{m}$ because the magnon frequency $\Omega$ is much smaller than the Fermi energy, whereas we assume that the polarization field $\bar{m}$ is comparable to the Fermi energy, as inferred from measurements [39]. We now show that $g^{(2)}(q)$ features a singularity at small momentum and frequency by examining two extreme regimes:
  1. (1)
    In the regime of $\nu\ll c_{s}|{\boldsymbol{q}}|$, $g_{l,x}^{(2)}(q)$ can be computed as follows. Taking $\nu=0$ in Eq. (75), we first massage the integrand to make it more symmetric—i.e., we shift the integration variable ${\boldsymbol{Q}}\rightarrow{\boldsymbol{Q}}-{\boldsymbol{q}}/2$ to obtain
    $g^{(2)}_{\rm l,x}(\nu=0,{\boldsymbol{q}})=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int_{\Omega,{\boldsymbol{Q}}}\frac{1}{\Omega^{2}+c_{s}^{2}({\boldsymbol{Q}}+\frac{{\boldsymbol{q}}}{2})^{2}}\frac{1}{\Omega^{2}+c_{s}^{2}({\boldsymbol{Q}}-\frac{{\boldsymbol{q}}}{2})^{2}}.$ (76)
    Taking ${\boldsymbol{q}}$ along the $x$ axis and denoting by $\theta$ the angle measured from ${\boldsymbol{q}}$, we then perform the Matsubara summation to obtain
    $g^{(2)}_{\rm l,x}(\nu=0,{\boldsymbol{q}})=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\frac{1}{2c_{s}^{3}|{\boldsymbol{q}}|}\int\frac{d|{\boldsymbol{Q}}|d\theta}{4\pi^{2}}\frac{1}{\cos\theta}\left(\frac{1}{\sqrt{{\boldsymbol{q}}^{2}+4{\boldsymbol{Q}}^{2}-4|{\boldsymbol{Q}}||{\boldsymbol{q}}|\cos\theta}}-\frac{1}{\sqrt{{\boldsymbol{q}}^{2}+4{\boldsymbol{Q}}^{2}+4|{\boldsymbol{Q}}||{\boldsymbol{q}}|\cos\theta}}\right).$ (77)
    The remaining integrals are complicated, but are well-behaved in the ${\boldsymbol{q}}\rightarrow 0$ limit of interest, where numerically they evaluate to exactly $1/4$. Thus we extract the leading behavior as
    $g^{(2)}_{\rm l,x}(\nu=0,{\boldsymbol{q}})\sim\frac{(V+J)\cos^{2}\theta_{0}}{32z_{s}D}\frac{1}{c_{s}|{\boldsymbol{q}}|},\quad\nu\ll c_{s}|{\boldsymbol{q}}|,$ (78)
    where we have used the definition of $c_{s}$ and $\kappa$ above to simplify the final result.
  2. (2)
    In the regime of $\nu\gg c_{s}|{\boldsymbol{q}}|$, we first perform the Matsubara frequency summation in Eq. (75) with ${\boldsymbol{q}}\rightarrow 0$, which yields
    $g^{(2)}_{\rm l,x}(\nu,{\boldsymbol{q}}=0)=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int\frac{d\Omega}{2\pi}\int\frac{d^{2}{\boldsymbol{Q}}}{4\pi^{2}}\frac{1}{(\Omega+\nu)^{2}+c_{s}^{2}{\boldsymbol{Q}}^{2}}\frac{1}{\Omega^{2}+c_{s}^{2}{\boldsymbol{Q}}^{2}}=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int\frac{d^{2}{\boldsymbol{Q}}}{4\pi^{2}}\frac{1}{c_{s}|{\boldsymbol{Q}}|}\frac{1}{\nu^{2}+4c^{2}_{s}{\boldsymbol{Q}}^{2}}.$ (79)
    In this limit the angular integral is trivial. Performing the radial momentum integral finally yields
    $g^{(2)}_{\rm l,x}(\nu,{\boldsymbol{q}}=0)=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\frac{1}{8c_{s}^{2}|\nu|}.$ (80)
    Replacing $c_{s}$ and $\kappa$ by its definitions above, we simplify this result to
    $g^{(2)}_{\rm l,x}(q)=\frac{(V+J)\cos^{2}\theta_{0}}{32z_{s}D}\frac{1}{|\nu|},\quad\nu\gg c_{s}|{\boldsymbol{q}}|.$ (81)
Therefore, the total pairing interaction arising from two-magnon processes (in the generic $\theta_{0}$ regime) is given by
$g_{\Phi}^{(2)}(q)=g^{(2)}_{\rm l}(q)+g^{(2)}_{\rm x}(q)\sim\begin{cases}\frac{I}{c_{s}|{\boldsymbol{q}}|},\quad\nu\ll c_{s}|{\boldsymbol{q}}|\\ \frac{I}{|\nu|},\quad\nu\gg c_{s}|{\boldsymbol{q}}|\end{cases},\quad I=\frac{V+J}{16z_{s}D}\cos^{2}\theta_{0},$ (82)
as quoted in Eq. (47) of the main text.
E.2 Ferromagnetic limit
Here we consider in more detail the ferromagnetic limit $\theta_{0}\rightarrow\pi/2$ where the divergent interaction in Eq. (50) vanishes. This behavior can be traced back to the fact that Goldstone magnons disperse quadratically in this limit. Our aim is to evaluate the pairing interaction generated by quadratically dispersing magnons with first-order scattering amplitudes given by Eq. (41). Evaluation of the crossed diagram represented in Fig. 4(b) yields
$g^{(2)}_{\Phi,\frac{\pi}{2}}(q)=\frac{1}{4\bar{m}^{2}}\int_{\Omega,{\boldsymbol{Q}}}\frac{1}{-i\gamma(\Omega+\nu)+D({\boldsymbol{Q}}+{\boldsymbol{q}})^{2}}\frac{1}{i\gamma\Omega+D{\boldsymbol{Q}}^{2}}.$ (83)
Performing first the Matsubara summation, we obtain
$g^{(2)}_{\Phi,\frac{\pi}{2}}(q)=\frac{1}{4\gamma\bar{m}^{2}}\int\frac{d^{2}{\boldsymbol{Q}}}{(2\pi)^{2}}\frac{1}{-i\gamma\nu+D{\boldsymbol{Q}}^{2}+D({\boldsymbol{Q}}+{\boldsymbol{q}})^{2}}.$ (84)
We once again shift the integration variable ${\boldsymbol{Q}}\rightarrow{\boldsymbol{Q}}-{\boldsymbol{q}}/2$ to render the expression more symmetric and, crucially, eliminate its angular dependence. Imposing a momentum cutoff $\Lambda$ and performing the remaining momentum integrals yields
@@ E.2 Ferromagnetic limit
$g^{(2)}_{\Phi,\frac{\pi}{2}}(q)=\frac{1}{8\pi\gamma\bar{m}^{2}}\int_{0}^{\Lambda}\frac{|{\boldsymbol{Q}}|d|{\boldsymbol{Q}}|}{-i\gamma\nu+2D\left({\boldsymbol{Q}}^{2}+{\boldsymbol{q}}^{2}/4\right)}=\frac{1}{32\pi\gamma\bar{m}^{2}D}\ln\frac{4\Lambda^{2}+D{\boldsymbol{q}}^{2}-2i\gamma\nu}{D{\boldsymbol{q}}^{2}-2i\gamma\nu}.$ (85)
In the zero frequency limit, $\nu=0$, we recover the result
$g^{(2)}_{\Phi,\frac{\pi}{2}}(\nu=0,{\boldsymbol{q}})=\frac{1}{32\pi\gamma\bar{m}^{2}D}\ln\frac{4\Lambda^{2}+{\boldsymbol{q}}^{2}}{{\boldsymbol{q}}^{2}}\sim\frac{1}{16\pi\gamma\bar{m}^{2}D}\ln\frac{2\Lambda}{|{\boldsymbol{q}}|},$ (86)
where we took the leading momentum dependence when $|{\boldsymbol{q}}|\ll\Lambda$ and discarded a constant $\sim\ln 2$. Finally, using the definition $\gamma^{-1}=2(V+J)\bar{m}$ we arrive at Eq. (53) from the main text.
The momentum-averaged pairing interaction on the Fermi surface is
$\overline{g^{(2)}_{\Phi,\frac{\pi}{2}}}(\nu=0)$ $=\int\frac{dq_{\parallel}}{2\pi k_{F}}g^{(2)}_{\Phi,\frac{\pi}{2}}(\nu=0,q_{\perp}=0,q_{\parallel})=\frac{q_{\rm max}}{32\pi^{2}\gamma\bar{m}^{2}Dk_{F}}\left(1+\ln\frac{\Lambda}{q_{\rm max}}\right).$ (87)
Importantly, this quantity is now finite—in stark contrast to the logarithmic low-frequency divergence obtained in the spin-canted case [cf. Eq. (50)].
E.3 Effect of gapping out the magnons with a Zeeman field
Next we revisit the analysis of scattering amplitudes in a situation where the Goldstone magnons are gapped out by applying a Zeeman field $B_{\parallel}$. The presence of $B_{\parallel}$ modifies the saddle-point analysis from Sec. II.3, e.g., by generically producing $O(B_{\parallel})$ shifts in the optimal canting angle and the net exchange field. (Non-analytic $B_{\parallel}$ dependence appears at the spin-canting transition where $\theta_{0}=0$, but we do not consider this special situation in what follows.) Such modifications in turn generate $O(B_{\parallel})$ corrections to the quantities $Q^{\phi,\theta}_{s,a}$ [Eq. (25)] that enter the magnon Green’s functions [Eq. (27)]. Upon including these corrections, the ladder/cross diagram amplitude from Eq. (75) becomes
$g^{(2)}_{\rm l,x}(q)=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int_{\Omega,{\boldsymbol{Q}}}\frac{1}{(\Omega+\nu)^{2}+c_{s}^{2}({\boldsymbol{Q}}+{\boldsymbol{q}})^{2}+a|B_{\parallel}|}\frac{1}{\Omega^{2}+c_{s}^{2}{\boldsymbol{Q}}^{2}+a|B_{\parallel}|}.$ (88)
The prefactor $a$ multiplying $|B_{\parallel}|$ is non-universal and can in principle be obtained from the modified saddle-point solution. For our purposes here, however, this general form suffices. By inspecting the integrand we see that the field-induced magnon gap for the Goldstone mode is $\Omega_{\rm mag}=\sqrt{a|B_{\parallel}|}$. We will focus on the limit ${\boldsymbol{q}}\rightarrow 0$ and ask how the low-frequency divergence is cut-off by gapping out the low-energy part of the magnon dispersion.
Evaluating the Matsubara summation first yields
$g^{(2)}_{\rm l,x}(\nu,{\boldsymbol{q}}=0)=\frac{\cos^{4}\theta_{0}}{16\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int\frac{d^{2}{\boldsymbol{Q}}}{(2\pi)^{2}}\frac{1}{\sqrt{c_{s}^{2}{\boldsymbol{Q}}^{2}+a|B_{\parallel}|}}\frac{1}{\nu^{2}+4(c^{2}_{s}{\boldsymbol{Q}}^{2}+a|B_{\parallel}|)},$ (89)
which can be compared to Eq. (79) above. Performing the momentum integrals then leads to
$g^{(2)}_{\rm l,x}(\nu,{\boldsymbol{q}}=0)=\frac{(V+J)\cos^{2}\theta_{0}}{32z_{s}D}\frac{1}{|\nu|}\left(1-\frac{2}{\pi}\arctan{\frac{2\sqrt{a|B_{\parallel}|}}{|\nu|}}\right),$ (90)
where we again replaced $c_{s}$ and $\kappa$ by their definition. Comparing now to Eq. (81), we conclude that the low-frequency divergence has been cut off by the inverse of the magnetic-field-induced magnon gap. This result can be seen explicitly by expanding the $\arctan$ function in the low-frequency limit to arrive at

### v3

@@ Appendix C Dispersion of magnons
C.1 Magnon spectrum: order of magnitude estimates and numerics
In this Section we provide estimates of the various scales involved in the magnon spectrum obtained from the poles of the magnon propagators $\mathcal{G}_{b\eta}(q)$ in Eq. (27) of the main text. For simplicity we first consider the critical point $\theta_{0}=0$, where the four magnon modes coalesce into only two solutions due to the restored U(1) spin rotation symmetry: a gapped ($+$) and a gapless, linearly dispersing (--) branch:
$\omega_{s\pm}({\boldsymbol{q}})=\omega_{a\pm}({\boldsymbol{q}})=\sqrt{2}\bar{m}\left[\alpha+2D(V+J){{\boldsymbol{q}}}^{2}\pm\sqrt{\alpha^{2}+4D(V+J){{\boldsymbol{q}}}^{2}}\right]^{1/2},$ (64)
where we defined the dimensionless parameter $\alpha=\frac{V+J}{V-J}$. In opposite limit of the SU(2) ferromagnet, with $\theta_{0}=\pi/2$, there are four non-degenerate magnon modes with dispersion
$\omega_{s\pm}({\boldsymbol{q}})$ $=\sqrt{2}\bar{m}\left[1+2D(V+J){{\boldsymbol{q}}}^{2}\pm\sqrt{1+4D(V+J){{\boldsymbol{q}}}^{2}}\right]^{1/2},$ (65)
$\omega_{a\pm}({\boldsymbol{q}})$ $=\sqrt{2}\bar{m}\left[\frac{V+3J}{V-J}+2D(V+J){{\boldsymbol{q}}}^{2}\pm\sqrt{\frac{V+7J}{V-J}+4D(V+J){{\boldsymbol{q}}}^{2}}\right]^{1/2}.$ (66)
Here the Goldstone ($s-$) mode is gapless (and quadratically dispersing), while the other three modes are gapped.
For rhombohedral multilayers (in particular following spin-orbit proximitized trilayer experiments [44]), we can estimate $J\sim 200$ meV nm2 and polarization densities (per valley) $n_{0}\sim 2\times 10^{11}$ cm2 relevant for the region where SC is observed. The strength of intra-valley interactions $V$ (which are momentum-independent within our microscopic treatment, Eq. (5)) can be roughly estimated from the value of the gate-screened Coulomb potential
$V(|{\boldsymbol{q}}|)=\frac{q_{\mathrm{e}}^{2}}{2\epsilon_{r}\epsilon_{0}|{\boldsymbol{q}}|}\tanh{(|{\boldsymbol{q}}|d)}$ (67)
at the Fermi momentum in the relevant low-density regime, $|{\boldsymbol{q}}|\sim k_{F}\sim\left(0.05-0.1\right)a_{0}^{-1}$. Here $q_{e}$ is the electron charge, $\epsilon_{0}$ the vaccum permittivity and $d$ denotes the gate distance, which is of order of a few tens of nanometers in typical experiments, such that $k_{F}d\gg 1$. The estimate for $V$ also depends on the amount of electronic screening in the system, which is difficult to evaluate accurately. In self-consistent Hartree-Fock treatments (e.g. [68, 52, 53, 54]) where screening is accounted for phenomenologically by treating the relative permittivity $\epsilon_{r}$ as a free parameter, a best fit to the experimentally observed phase boundaries [3, 5, 40, 44] is obtained in the range $\epsilon_{r}\sim 20-30$. With these various uncertainties in mind, we estimate $V(|{\boldsymbol{q}}|\sim k_{F})\sim 1-3$ eV nm2. The corresponding ratio of intra-valley to inter-valley interactions in our theory should then be $V/J\sim 5-15$.
Using these values, the induced mass gap $\bar{m}=n_{0}(J+V)\sim 2-6$ meV compares reasonably to Hartree-Fock simulations and is of the same order (although a few times larger) than the experimentally extracted peak splittings using STM in single-gated rhombohedral trilayer devices [74]. As a further consistency check, such values of $V$ (which describe a local interaction term in our theory), are of the correct order for the onset of Stoner transitions, $VD_{F}\approx 1$, considering typical density of states $D_{F}\sim 1$ eV-1 nm-2 in rhombohedral graphene systems near charge neutrality under a large applied displacement field.
The parameter $D$ depends sensitively on the band dispersion, including trigonal warping and corrections due to interactions, as well as the electronic density of interest. While we do not attempt a careful, quantitative calculation here (which should be carried out self-consistently), we can estimate the orders of magnitude involved. The susceptibility $\chi_{1}(q)$ has units of density of states, and we therefore take $\chi\sim 1$ eV-1 nm-2 at $|{\boldsymbol{q}}|\sim k_{F}\sim 0.1a_{0}^{-1}\sim 0.4$ nm-1 such that $D\sim 6$ eV-1.
We plot the magnon spectrum in Fig. 3 for $D=1$ and $10$ eV-1, and two values of $V/J=5,10$, both in the ranges justified above. The separation of energy scales between the high-energy and the soft magnon modes is quite large for momentum transfers of the order $k_{0}a_{0}\sim 0.01-0.05$, which are expected to be most important for interband pairing that scatter electrons between majority and minority Fermi surfaces. (Note that $k_{0}$ denotes the momentum transfer between the majority and minority Fermi pockets, which can be much smaller than $k_{F}$ especially in the presence of trigonal warping).
Figure 3: Magnon spectrum evaluated numerically from Eqs. (64)–(66). We consider the critical point, $\theta_{0}=0$, for (a) $V/J=5$ and (b) $V/J=10$, and (c) the SU(2)-symmetric ferromagnet, $\theta_{0}=\pi/2$, for $V/J=10$. The four modes are color-coded as follows: (blue = $s-$, Goldstone mode), (green = $a-$, near-critical soft mode), (orange = $s+$ and red = $a+$, high-energy modes). The characteristic scale for the minimal transferred momentum between majority and minority pockets, $k_{0}$, is highlighted in grey, and the value of $q_{\rm max}$ that characterizes the linear regime where the Goldstone-mode dispersion Eq. (29) is valid (at the critical point) is denoted by black vertical lines. Inspired by rhombohedral graphene experiments, in particular spin-orbit proximitized trilayers [44], we use the Hund’s coupling scale $J=200$ meV nm2 and a polarization density (per valley) $n_{0}=2\times 10^{11}$ cm-2 throughout, for two values for $D=1$ eV-1 (solid lines) and $10$ eV-1 (dashed lines) . The gaps to the two high-energy ($+$) modes are set primarily by the saddle-point value of the order parameter, $\bar{m}=n_{0}(V+J)$.
Appendix D Derivation of magnon-mediated interaction
The spin fluctuation we considered in the last section mediates an electron-electron interaction of the following form:
$\mathcal{S}_{I}$ $=\sum_{kpq}\sum_{\xi_{1},\xi_{2}\in\{K,K^{\prime}\}}\sum_{\alpha\alpha^{\prime}\beta\beta^{\prime}=\pm}g_{\xi_{1};\xi_{2}}^{\alpha^{\prime}\alpha;\beta^{\prime}\beta}(q)\overline{\Psi}_{\xi_{1},k+q,\alpha^{\prime}}\overline{\Psi}_{\xi_{2},p-q,\beta^{\prime}}\Psi_{\xi_{2},p,\beta}\Psi_{\xi_{1},k,\alpha},$ (68)
where the scattering amplitude matrix $g$ is given by
$g_{\xi_{1},\xi_{2}}^{\alpha^{\prime}\alpha,\beta^{\prime}\beta}(q)$ $=-\sum_{\mu\lambda}\mathcal{M}_{\xi_{1}\xi_{2}}^{\mu\lambda}(q)\sigma^{\mu}_{\alpha^{\prime}\alpha}\sigma^{\lambda}_{\beta^{\prime}\beta},$ (69)
where $\mu,\lambda\in\{\theta,\phi\}$, $\sigma^{\theta}=\left(\begin{matrix}0&1\\ 1&0\end{matrix}\right)$, $\sigma^{\phi}=\left(\begin{matrix}0&-i\\ i&0\end{matrix}\right)$, whereas $\mathcal{M}_{\xi_{1}\xi_{2}}$ is defined as
$\mathcal{M}_{\xi_{1}\xi_{2}}$ $=\left\langle\left(\begin{matrix}\delta m^{\theta}_{\xi_{1}}\delta m^{\theta}_{\xi_{2}}&\delta m^{\theta}_{\xi_{1}}\delta m^{\phi}_{\xi_{2}}\\ \delta m^{\phi}_{\xi_{1}}\delta m^{\theta}_{\xi_{2}}&\delta m^{\phi}_{\xi_{1}}\delta m^{\phi}_{\xi_{2}}\end{matrix}\right)\right\rangle$ (70)
Using definition of symmetric and antisymmetric modes in Eq. (23), we express $\mathcal{M}$ in terms of symmetric and antisymmetric modes’ contributions:
$\mathcal{M}_{\xi_{1}\xi_{2}}$ $=\frac{1}{2}\langle\delta M_{s}\delta M_{s}\rangle+\frac{1}{2}\tau_{\xi_{1}}\tau_{\xi_{2}}\langle\delta M_{a}\delta M_{a}\rangle$ (71)
Both terms can be read off from Eq. (24) by taking the inverse of the 2-by-2 matrix in Eq. (24):
@@ Appendix D Derivation of magnon-mediated interaction
$\mathcal{M}_{\xi_{1}\xi_{2}}=\frac{1}{2}\frac{1}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}\left(\begin{matrix}Q^{\phi}_{s}&-\gamma\nu_{n}\\ +\gamma\nu_{n}&Q^{\theta}_{s}\end{matrix}\right)+\frac{1}{2}\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}}\left(\begin{matrix}Q^{\phi}_{a}&-\gamma\nu_{n}\\ +\gamma\nu_{n}&Q^{\theta}_{a}\end{matrix}\right)$ (72)
where $\tau_{\xi}=\pm 1$ for $\xi=K,K^{\prime}$. So $\tau_{\xi_{1}}\tau_{\xi_{2}}=\pm 1$ represents whether $\xi_{1}$ and $\xi_{2}$ are the same valley indices. This factor enters the second term because the second term represents the contribution of the antisymmetric magnon modes which feature opposite amplitudes in valleys $K$ and $K^{\prime}$. Next, we explicitly compute all the magnon-mediated scattering amplitudes
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu_{n})+(-i)(\gamma\nu_{n})+(-i)^{2}Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu_{n})+(-i)(\gamma\nu_{n})+(-i)^{2}Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}-Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}-Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$ (73)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu_{n})+i(\gamma\nu_{n})+i(-i)Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu_{n})+i(\gamma\nu_{n})+i(-i)Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}+Q^{\theta}_{s}+2i\gamma\nu_{n}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+Q^{\theta}_{a}+2i\gamma\nu_{n}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}}$ (74)
where
$Q^{\theta}_{s}=2\tilde{J}\cos^{2}\theta_{0}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{s}=\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},$ (75)
$Q^{\theta}_{a}=2\tilde{J}\sin^{2}\theta_{0}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{a}=2\tilde{J}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2}.$ (76)
As a consistency check, we look at the case of canting critical point where $\theta_{0}=0$. Eqs. (73) and (74) show the intravalley $g^{-+,-+}_{\xi,\xi}$ and intervalley $g^{-+,+-}_{K,K^{\prime}}$ amplitudes vanish, in line with expectations from symmetry: At $\theta_{0}=0$, the system preserves global U(1) spin rotation symmetry around the $z$ axis, which in our rotated basis enforces conservation of $S^{z}_{\rm tot}=\sum_{\xi}\int_{k}\overline{\Psi}^{\dagger}_{\xi,k}\tau_{\xi}\sigma^{z}\Psi_{\xi,k}$. Intravalley $g^{-+,-+}_{\xi,\xi}$ amplitudes encode processes that change $S^{z}_{\rm tot}$ by flipping two spins in a given valley; intervalley $g^{-+,+-}_{K,K^{\prime}}$ amplitudes similarly violate this conservation law.
Next, plugging Eq. (75) and Eq. (76) into Eq. (73) and Eq. (74), and dropping higher-order terms in $q$ and $\nu$ yields